OpenStudy (anonymous):

using the limit definition of the derivative find f'(x). Let f(x) = x^2 -4

OpenStudy (therealmeeeee):

f(x) = x^3 - 7x + 5 f(x + h) = (x + h)^3 - 7 * (x + h) + 5 f(x + h) - f(x) => (x + h)^3 - 7 * (x + h) + 5 - x^3 + 7x - 5 => x^3 + 3x^2 * h + 3x * h^2 + h^3 - 7x - 7h - x^3 + 7x => 3x^2 * h + 3x * h^2 + h^3 - 7h Divide by h 3x^2 + 3xh + h^2 - 7 Let h go to 0 3x^2 - 7 x = 5 3 * 5^2 - 7 => 3 * 25 - 7 => 75 - 7 => 68 So we need a line with a slope of 68 that passes through (5 , 95) y - 95 = 68 * (x - 5) y - 95 = 68x - 340 y = 68x - 245 f(x) = x^3 - 7x + 5 f(x + h) = (x + h)^3 - 7 * (x + h) + 5 f(x + h) - f(x) => (x + h)^3 - 7 * (x + h) + 5 - x^3 + 7x - 5 => x^3 + 3x^2 * h + 3x * h^2 + h^3 - 7x - 7h - x^3 + 7x => 3x^2 * h + 3x * h^2 + h^3 - 7h Divide by h 3x^2 + 3xh + h^2 - 7 Let h go to 0 3x^2 - 7 x = 5 3 * 5^2 - 7 => 3 * 25 - 7 => 75 - 7 => 68 So we need a line with a slope of 68 that passes through (5 , 95) y - 95 = 68 * (x - 5) y - 95 = 68x - 340 y = 68x - 245

OpenStudy (anonymous):

what?

OpenStudy (therealmeeeee):

this your answer

OpenStudy (therealmeeeee):

medals now for me helping u

OpenStudy (anonymous):

That's not the answer. I don't have a 7 anywhere in my problem.

OpenStudy (dumbcow):

@TheRealMeeeee stop copy/pasting random solutions that do not apply

OpenStudy (therealmeeeee):

im not so shut up

OpenStudy (therealmeeeee):

i wrote that myself

OpenStudy (therealmeeeee):

stupid lil boy

OpenStudy (anonymous):

Well, you're wrong

OpenStudy (therealmeeeee):

who

OpenStudy (therealmeeeee):

http://openstudy.com/updates/529ce572e4b0b97c585d97da

OpenStudy (dumbcow):

def of derivative \[\lim_{h \rightarrow 0} \frac{f(x+h) -f(x)}{h}\] f(x) = x^2 -4 f(x+h) = (x+h)^2 -4 = x^2+2hx +h^2 -4 \[\lim_{h \rightarrow 0}\frac{(x^{2}+2xh+h^{2}-4) - (x^{2}-4)}{h} = \frac{2xh +h^{2}}{h}\] now just factor out an "h" and see that it cancels with "h" on bottom , then apply limit

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

yw later you will learn a shortcut for finding these derivatives and you wont have to use the long limit process :)